Draw the shear and moment diagrams. 6 kN 6 kN/m B -3 m -1.5 m 1.5 m
Answer #1

6 KM TMm Ah Dhim - 3mpx lith S70 + 1B RAARB = 8+£x3x8| Ratre = ls-0 1125 Bip Moment @A » Rexo =(x +3 + . RB = 7.5 KN mp K - +